OpenStudy (anonymous):

What's log base 4 to 9 ? Evaluate. Show all your steps.

OpenStudy (king):

\[\log_{4} 9\]

OpenStudy (anonymous):

yup, thats the one

OpenStudy (king):

is \[ \log_{2} 3\]

OpenStudy (king):

see 9=3^2 and 4=2^2 so 2/2*log base 2 to 3 that is log base 2 to 3

OpenStudy (anonymous):

Write 4 as 2^2 and 9 as 3^2 and then use the power law to simplify.

OpenStudy (anonymous):

power law?

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

okay, got it. THANKS GUYS!